Slove the system of equations by the addition method
Find x & y
-3x = -5y-12
6x+3y = -15

Answer by checkley77(12569)

-3x+5y=-12 multiply by 2 & add.
6x+3y=-15
-6x+10y=-24
--------------------
13y=-39
y=-39/13
y=-3 ans.
-3x+5*-3=-12
-3x-15=-12
-3x=-12+15
-3x=3
x=3/-3
x=-1